OpenStudy (anonymous):

At what point does the line given by the following equation cross the x-axis? y = -5/7x + 5

OpenStudy (anonymous):

when y=0 0=-5/7x+5 -5/7x=-5 -35x=-5 x=-5/-35 x=1/7

OpenStudy (anonymous):

okay thanks you

OpenStudy (anonymous):

Hamilton received an 81 on his algebra exam. The professor is allowing students to take a 5-question bonus quiz to improve their algebra exam score. Each question answered correctly on the bonus quiz will add 3 points to the exam score. Hamilton needs to receive at least a 90 on his exam score in order to make an "A" for the semester. If x represents the number of questions answered correctly on the bonus quiz, which of the following inequalities symbolizes this situation? A. 5x + 81 < 90 B. 3x + 81 > 90 C. 3x + 81 < 90 D. 5x + 81 > 90

OpenStudy (anonymous):

@captaincook

OpenStudy (anonymous):

B